Fermi National Laboratory
 

Accelerator Update for December 13, 2006 - December 15, 2006

 

Wednesday December 13

The day shift began with Operations monitoring store 5119 (37.46E30) and a Recycler stash (216.7E10), with the Antiproton Source (Pbar) stacking (59.6E10), with Linac, Booster, Main Injector (MI), and Recycler operating normally, with MiniBooNE off due to LCW trouble, with NuMI taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

The Recycler completed a stash of antiprotons at 8:40 AM.

Operations terminated store 5119 at 9:01 AM.

Operations prepared Pbar and the Switchyard (SY) enclosure B for access at 9:01 AM.

Experts began the MI abort kicker repairs at 9:04 AM.

Operations started running beam to the Booster absorber at 9:18 AM.

Experts began work on the Booster RF system at 10:05 AM.

At 10:42 AM, Operations reported that they had prepared the following areas for access: Tevatron (TeV) sectors A-E, Transfer Hall, F-sector, Pbar, MI-12B, SY enclosures B, C, D, & E, D0 and CDF collision halls, and the Meson M02-M05 tunnels.

CDF came out of access 3:46 PM.

A MI kicker expert reported at 5 PM that they had completed the kicker work. Both kickers had been tested to 34 kV. He said the main source of the problem was a failed power supply.

The MiniBooNE LCW work was completed. Operations established beam to MiniBooNE at 5 PM.

Pbar came out of access at 6:25 PM.

TeV came out of access at 6:29 PM.

Operations established beam to NuMI at 7:11 PM.

Pbar began stacking at 7:30 PM.

Pbar experts began a study period at 8:05 PM.

Operations put the TeV through a dry squeeze at 10:44 PM.

Operations put the TeV through a wet squeeze at 11:43 PM.

Thursday December 14

The midnight shift began with Operations monitoring a stash (259.9E10), with Pbar stacking (49.9E10), and with NuMI and MiniBooNE taking beam.

Operations put the TeV into shot setup at 12:17 AM.

The NuMI horn tripped off at 12:38 AM. It reset okay.

Operations established store 5123 at 2:22 AM with an initial luminosity of 197.2E30.

Tevatron Electron Lens (TEL) experts began their TEL2 studies at 2:27 AM. The studies lasted about two hours.

At 2:45 AM, D0 reported that they were having network problems, which were affecting luminosity readbacks.

The Recycler completed a stash at 2:55 AM.

At 5:34 AM, the I- Source reference tripped and wouldn’t reset. Operators swapped out the reference module with no effect and then reinstalled the original module. Normal operations resumed at 6:03 AM.

The Recycler completed a stash at 6:19 AM.

At 9:13 AM, the TeV suffered a sector B0 quench. The store was lost. The current for a sector B0 quadrupole (B0Q6) had dropped to zero. Experts investigated.

Operations prepared CDF for access at 11:03 AM.

The Recycler completed a stash 12:09 PM.

Operators began preparing SY and the Meson area for the T959 experiment at 2:22 PM.

At 3:06 PM, Operations reported a LCW leak in Pbar’s PreVault area (AP1).

At 3:12 PM, Operations reported that while they were preparing the TeV for a dry squeeze, the TeV quenched at sectors A4, BA, BB, and BD.

Operations prepared the AP1 PreVault area for access at 3:14 PM.

Operations reported at 4:35 PM that a Pbar magnet (EB6) had a LCW hose break. Operators had the water squeegeed from the PreVault area by 5:32 PM.

Pbar resumed stacking by 5:47 PM.

Operations had D0 prepared for a forty-five minute access at 6:16 PM.

TeV experts began a study period at 8:03 PM. The study lasted two hours.

Operations put the TeV into shot setup at 10:52 PM.

Friday December 15

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (99.2E10), and with MiniBooNE and NuMI taking beam.

Operations established store 5127 at 1:01 AM with an initial luminosity of 190.91E30.

The Recycler completed it last of three stashes on the midnight shift at 4:22 AM.

At 4:22 AM, the NuMI beam permit tripped due a kicker temperature spike for the first of four times in about three hours. It reset each time.

At 7:16 AM, Operations reported that they could not keep the NuMI beam permit up due to the kicker temperature trips. Operations contacted an expert who commented that the kicker level was okay, but the cooling system level was a little low. The expert will fill this when he gets a chance. With some expert help, operators made some adjustments to the alarm limit and reset the beam permit.

The Plans for Friday and the Weekend

The plans for today and this weekend are to stack, stash, and store. Experts and Operations will attempt to start sending beam to the Meson MCenter T959 experiment this afternoon. The Run Coordinator will allow studies that don’t interfere with the store, the stash, or the stack.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 12/15/2006    email Fermilab